# Carré du champ flow matching: better quality-generalisation tradeoff in generative models

### v1

@@ 4.3 Limits of explicit geometric regularisation
As before, we found that FM samples were of consistently high quality, and while CDC-FM could attain a comparable quality, it required an increasingly larger number of training points as the dimension increased (Fig. 5a). Yet, strikingly, FM memorised almost all of the dataset when the dimension was high enough (Fig. 5b). By contrast, while memorisation in FM and CDC-FM was comparable in dimensions one or two, CDC-FM memorisation decreased with dimension and remained low (Fig. 5b). CDC-FM also attained generally higher generalisation (Fig. 5c). Taken together, these experiments show that, unlike FM, CDC-FM prevents memorisation and facilitates generalisation irrespective of dimension, an effect that remained robust when adding small Gaussian noise to the data (Fig. A4). Yet, as the dimension increases, the desirable sample quality may only be achievable with a sufficient amount of data, likely due to the kknn graph construction of the kernel.
@@ 5 Related work
Our approach is motivated by geometric approaches of generative modelling under the so-called manifold hypothesis, surveyed in loaiza-ganem_deep_2024, which presumes that high-dimensional data often concentrates near a lower-dimensional manifold. Generative models have also been used to estimate intrinsic geometric quantities, e.g., manifold dimension (stanczuk_your_2023). Conversely, related line of work constrains generative models to predefined manifolds (chen_flow_2024; huang_riemannian_2022; mathieu_riemannian_2020; bortoli_convergence_2023), or to manifolds that are learnt from data (kapusniak2024metric; peach_implicit_2024; gosztolai2025). Our work also uses manifold tangent space estimates, but differs by focusing on modifying the probability path with an anisotropic diffusive term aligned to estimated tangents, rather than constraining generation to a pre-specified manifold.
@@
Memorisation and generalisation have been studied in the context of diffusion models, where empirical evidence showed a strong dependence on training set size and architecture (yoon2023diffusion; gu_memorization_2023). From a geometric standpoint, ross2025a argues that memorisation occurs when the learned manifold’s dimension is too low due to either overfitting-driven memorisation, when tangent directions are not fully captured, or data-driven memorisation, when the underlying data manifold itself is degenerate. In this regard, achilli_losing_2024 observed failures to capture tangent space dimensions and ventura_manifolds_2025 studied the dynamical regimes of diffusion models from a geometric perspective. On the theoretical side, bortoli_convergence_2023 studied the convergence of denoising diffusion models under manifold assumptions. While FM and diffusion share strong formal analogies — indeed, FM unifies score-based and diffusion training — there has been much less work on memorisation in FM. bertrand2025closed shows that the optimal FM vector field memorises and argues that generalisation does not come from the stochasticity of the FM objective. Our results thus complement and extend the existing understanding of memorisation phenomena in generative modelling.
@@ 6 Discussion
One limitation of our method stems from the use of the manifold hypothesis. We found that as the manifold dimension rises, our method needs exponentially more samples to maintain accuracy, due to the need to estimate tangent spaces from local neighbourhoods. Further, for non-geometric data, the benefit of geometric regularisation diminishes, on average, as the training set size increases, because the neural network architectures and the loss function already confer inductive biases.
Yet, we show that sufficient implicit regularisation can depend heavily on the local complexity and sparsity of the data, while the use of CDC-FM does not degrade performance. Thus, we expect that even in overall high data settings, local memorisation is likely a common occurrence, driven by local sparsity patterns (Skrinjar). In these scenarios, we expect that CDC-FM can provide a robust tool to reduce or eliminate memorisation and improve generalisation. Thus, our approach is not a competitor to FM but a plug‑in regulariser that can be scheduled, adapted, or even learned, opening a path to geometry‑aware flow-based generative models with stronger guarantees, better sample efficiency, and improved robustness to privacy risks.
@@ 9 Reproducibility statement
All details necessary to reproduce the experiments are present in the paper. We commit to publishing an open-source implementation of our code and guidelines to reproduce each experiment upon publication.
@@ Appendix B Carré du champ flow matching generates optimal transport flow paths
because, if $X\sim\tilde{p}_{t}(\cdot|X_{1})$, then X=ta+(1t)bX=ta+(1-t)b where $a\sim\mathcal{N}(0,\mathbf{I})$ and $b\sim\mathcal{N}(x_{1},\widehat{\mathbf{\Gamma}}(x_{1}))$ are independently sampled. but the optimal probability path between $\mathcal{N}(0,\mathbf{I})$ and $\mathcal{N}(x_{1},\widehat{\mathbf{\Gamma}}(x_{1}))$ is given by the displacement interpolant
@@
$=\mathcal{N}\big(tX_{1},(1-t)^{2}\mathbf{I}+t^{2}\widehat{\mathbf{\Gamma}}(X_{1})^{1/2}+2t(1-t)\widehat{\mathbf{\Gamma}}(X_{1})^{1/2}\big),$
@@
Appendix C Interpretation of CDC-FM as an anisotropic diffusive regularisation
@@
$\dot{\mathbf{\Sigma}}_{t}(x_{1})=\left[(1-t)\mathbf{I}+t\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}\right](\widehat{\mathbf{\Gamma}}^{1/2}-\mathbf{I}).$
@@
Appendix D Pseudo code
Algorithm 1 Flow Matching
1:Input: Samples $\mathcal{D}=\{x^{(i)}\}^{N}_{i=1}\subset\mathbb{R}^{d}$ from $\nu(x)$ parameterized vector field $v^{\theta}(x,t)$.
2:
3:
5:Sample $x_{1}\leftarrow x^{(i)}\sim\nu(x)$
6:Sample $t\sim\text{Uniform}[0,1]$
7:Sample $x_{0}\sim\mathcal{N}\left(0,\mathbf{I}\right)$
8:Define
9:
10:Define the target conditional vector field at
$v_{t}(x_{t}\mid x_{1})\leftarrow x_{1}-\left(1-\sigma_{\min}\right)x_{0}$
11:
12:Compute the loss
13:Compute the gradient $\nabla_{\theta}\mathcal{L}(\theta)$
14:return $\nabla_{\theta}\mathcal{L}(\theta)$
Algorithm 2 CDC Flow Matching
1:Input: Samples $\mathcal{D}=\{x^{(i)}\}^{N}_{i=1}\subset\mathbb{R}^{d}$ from $\nu(x)$ together with their local covariance square-root $\widehat{\mathbf{\Gamma}}(x^{(i)})^{0.5}$; parameterized vector field $v^{\theta}(x,t)$.
3:Sample $x_{1}\leftarrow x^{(i)}\sim\nu(x)$
4:Sample $t\sim\text{Uniform}[0,1]$
5:Sample $x_{0}\sim\mathcal{N}\left(0,\mathbf{I}\right)$
6:Define
7:Define the target conditional vector field
8:Compute the loss
9:Compute the gradient $\nabla_{\theta}\mathcal{L}(\theta)$
10:return $\nabla_{\theta}\mathcal{L}(\theta)$
@@ Appendix E Optimal estimation of the carré du champ matrix
For $\Sigma\succ 0$, the log-likelihood of a sample $Y\\ N(m,\mathbf{\Sigma})$ is
@@ Appendix F Computational complexity and runtimes
Table A2: Preprocessing empirical runtimes. Runtimes are reported in seconds for CIFAR-10 subsets of varying sizes and were ran on NVIDIA A10s.
@@ Appendix H Supplementary tables
Appendix I Supplementary figures

### v2

@@ 4.3 Limits of explicit geometric regularisation
As before, we found that FM samples were of consistently high quality, and while CDC-FM could attain a comparable quality, it required an increasingly larger number of training points as the dimension increased (Fig. 5a). Yet, strikingly, this increased quality by FM was a result of memorising almost all of the dataset when the dimension was high enough (Fig. 5b). This shows that memorisation is more likely to occur for higher dimensions for a fixed data size. By contrast, while memorisation in FM and CDC-FM was comparable in dimensions one or two, CDC-FM memorisation decreased with dimension and remained low (Fig. 5b). CDC-FM also attained generally higher generalisation (Fig. 5c). Taken together, these experiments show that, unlike FM, CDC-FM prevents memorisation and facilitates generalisation irrespective of dimension, an effect that remained robust when adding small Gaussian noise to the data (Fig. A4). Yet, as the dimension increases, the desirable sample quality may only be achievable with a sufficient amount of data, likely due to the kknn graph construction of the kernel.
@@
Addressing the curse of dimension through latent diffusion: Celeba-HQ.
Table 2: FID and NLL per epoch for FM vs. CDC-FM in latent space. CelebA-HQ subset of size $1000$.
Epoch FID NLL
FM CDC-FM FM CDC-FM
1000 15.60 12.72 6.80 7.18
2000 13.51 13.42 6.78 6.83
3000 13.56 10.55 6.80 6.68
4000 13.82 11.70 6.69 6.53
5000 13.10 10.85 6.68 6.48
We then trained FM and CDC-FM in latent space, and decoded the output back to pixel space. To test whether our findings extend to such a setting, and following the setup of dao2023flow, we evaluate FM and CDC-FM in the latent space of the Stable Diffusion VAE (rombach2022high). We train both models on a subset of $1000$ Celeba-HQ images, and report FID and NLL across epochs in Table 4.3. Despite the dimensionality reduction to size $32\times 32\times 4$, we observe that the CDCFM regularisation improves both quality and generalisation after 33k epochs once both model performances stabilise.
@@ 5 Related work
Our approach is motivated by geometric approaches of generative modelling under the so-called manifold hypothesis, surveyed in loaiza-ganem_deep_2024, which presumes that high-dimensional data often have low-dimensional structure. Generative models have also been used to estimate intrinsic geometric quantities, like dimension (stanczuk_your_2023). Conversely, related line of work constrains generative models to predefined manifolds (chen_flow_2024; huang_riemannian_2022; mathieu_riemannian_2020; bortoli_convergence_2023), or to manifolds that are learnt from data (kapusniak2024metric; peach_implicit_2024; gosztolai2025). Our use of the carré du champ is based on diffusion geometry (jones2024diffusion), which generalises classical Riemannian geometry to complex geometries that do not meet the strict definition of a manifold. Rather than using the geometry as a constraint, we use it as a form of model regularisation by modifying the probability path with an anisotropic diffusive term aligned to the data geometry.
@@
Memorisation and generalisation have been studied in the context of diffusion models, where empirical evidence showed a strong dependence on training set size and architecture (yoon2023diffusion; gu_memorization_2023). From a geometric standpoint, ross2025a argues that memorisation occurs when the learnt manifold’s dimension is too low due to either overfitting-driven memorisation, when tangent directions are not fully captured, or data-driven memorisation, when the underlying data manifold itself is degenerate. In this regard, achilli_losing_2024 observed failures to capture tangent space dimensions and ventura_manifolds_2025 studied the dynamical regimes of diffusion models from a geometric perspective. On the theoretical side, bortoli_convergence_2023 studied the convergence of denoising diffusion models under manifold assumptions. While FM and diffusion share strong formal analogies — indeed, FM unifies score-based and diffusion training — there has been much less work on memorisation in FM. bertrand2025closed shows that the optimal FM vector field memorises and argues that generalisation does not come from the stochasticity of the FM objective. Our results thus complement and extend the existing understanding of memorisation phenomena in generative modelling.
@@ 6 Discussion
Our model uses a form of the manifold hypothesis based on local heat kernel approximations. This is weaker than the typical Riemannian approximation, as it allows for the intrinsic dimension to vary and is robust to noise. Yet, we found that as the manifold dimension rises, our method needed exponentially more samples to maintain accuracy, due to the need to estimate tangent spaces from local neighbourhoods. Further, for non-geometric data, the benefit of geometric regularisation diminished, on average, as the training set size increased, because the neural network architectures and the loss function already confer inductive biases. Yet, we showed that architecture-induced regularisation can depend heavily on the local complexity and sparsity of the data, while the use of CDC-FM remained robust and resulted in better generalisation and, in some cases, also higher quality in diverse data settings than FM.
Overall, we expect that even in overall high data settings, local memorisation is likely a common occurrence, driven by local sparsity patterns (Skrinjar). In these scenarios, we expect that CDC-FM can provide a robust tool to reduce or eliminate memorisation and improve generalisation. Thus, our approach is not a competitor to FM but a plug‑in regulariser that can be scheduled, adapted, or even learned and can be readily used in conjunction with existing latent space pipelines. While we used our geometric framework to generalise FM, being one of the most broadly adopted generative models, future work could explore the effect of our geometric regularisation on other generative models and stochastic regularisation strategies. Our work thus opens a path to geometry‑aware flow-based generative models with stronger guarantees, better sample efficiency, and improved robustness to privacy risks.
@@ 9 Reproducibility statement
All details necessary to reproduce the experiments are present in the paper. Code is available at https://github.com/Dynamics-of-Neural-Systems-Lab/cdc-fm.
@@ Appendix B Carré du champ flow matching generates optimal transport flow paths
because, if $X\sim\tilde{p}_{t}(\cdot|X_{1})$, then X=ta+(1t)bX=ta+(1-t)b where $a\sim\mathcal{N}(0,\mathbf{I})$ and $b\sim\mathcal{N}(x_{1},\widehat{\mathbf{\Gamma}}(x_{1}))$ are independently sampled, but the optimal probability path between $\mathcal{N}(0,\mathbf{I})$ and $\mathcal{N}(x_{1},\widehat{\mathbf{\Gamma}}(x_{1}))$ is given by the displacement interpolant
@@
$=\mathcal{N}\big(tX_{1},(1-t)^{2}\mathbf{I}+t^{2}\widehat{\mathbf{\Gamma}}(X_{1})+2t(1-t)\widehat{\mathbf{\Gamma}}(X_{1})^{1/2}\big),$
@@
Appendix C Pseudo code
Algorithm 1 Flow Matching
1:Input: Samples $\mathcal{D}=\{x^{(i)}\}^{N}_{i=1}\subset\mathbb{R}^{d}$ from $\nu(x)$ parameterised vector field $v^{\theta}(x,t)$.
2:
3:
5:Sample $x_{1}\leftarrow x^{(i)}\sim\nu(x)$
6:Sample $t\sim\text{Uniform}[0,1]$
7:Sample $x_{0}\sim\mathcal{N}\left(0,\mathbf{I}\right)$
8:Define
9:
10:Define the target conditional vector field at
$v_{t}(x_{t}\mid x_{1})\leftarrow x_{1}-\left(1-\sigma_{\min}\right)x_{0}$
11:
12:Compute the loss
13:Compute the gradient $\nabla_{\theta}\mathcal{L}(\theta)$
14:return $\nabla_{\theta}\mathcal{L}(\theta)$
Algorithm 2 CDC Flow Matching
1:Input: Samples $\mathcal{D}=\{x^{(i)}\}^{N}_{i=1}\subset\mathbb{R}^{d}$ from $\nu(x)$ together with their local covariance square-root $\widehat{\mathbf{\Gamma}}(x^{(i)})^{0.5}$; parameterised vector field $v^{\theta}(x,t)$.
3:Sample $x_{1}\leftarrow x^{(i)}\sim\nu(x)$
4:Sample $t\sim\text{Uniform}[0,1]$
5:Sample $x_{0}\sim\mathcal{N}\left(0,\mathbf{I}\right)$
6:Define
7:Define the target conditional vector field
8:Compute the loss
9:Compute the gradient $\nabla_{\theta}\mathcal{L}(\theta)$
10:return $\nabla_{\theta}\mathcal{L}(\theta)$
Appendix D Interpretation of CDC-FM as an anisotropic diffusive regularisation
@@
$\dot{\mathbf{\Sigma}}_{t}(x_{1})=\left[(1-t)\mathbf{I}+t\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}\right](\widehat{\mathbf{\Gamma}}(x_{1})^{1/2}-\mathbf{I}).$
@@ Appendix E Optimal estimation of the carré du champ matrix
For $\Sigma\succ 0$, the log-likelihood of a sample $Y\sim\mathcal{N}(m,\mathbf{\Sigma})$ is
@@ Appendix F Computational complexity and runtimes
Table A2: Preprocessing empirical runtimes. Runtimes are reported in seconds for CIFAR-10 subsets of varying sizes and were run on NVIDIA A10s.
@@ Appendix G Experiment details
Celeba-HQ experiment.
For the Celeba-HQ, we followed the setup in dao2023flow with the exception that we did not use the standard flip transform for data augmentation to focus on the fundamental differences between FM and CDC-FM losses. We use a UNet (dhariwal2021diffusion). For NLL we report bits per dimension of a test set in latent space. For FID computation we use the inception v3 embeddings (szegedy2016rethinking). The CDC-FM-specific parameters were $k=256$, $k_{bw}=8$, $d_{cdc}=4$, $\gamma=0.5$.
@@ Appendix H Supplementary tables
Appendix I Negative Log-Likelihood Computation
To evaluate the generalisation capability of our models, we compute the negative log-likelihood (NLL) on held-out test data. Recall that the flow matching model defines a vector field $u_{t}(x)$ that generates a probability path $p_{t}(x)$ transforming a source distribution $p_{0}$ (typically $\mathcal{N}(0,\mathbf{I})$) to the model distribution $p_{1}\approx\nu$. The log-density of a test point $x$ under the model is computed using the instantaneous change of variables formula for Continuous Normalising Flows (chen_neural_2019):
$\log p_{1}(x)=\log p_{0}(\psi_{1}^{-1}(x))-\int_{0}^{1}\nabla\cdot u_{t}(\psi_{t}(\psi_{1}^{-1}(x)))\,dt,$ (34)
where $\psi_{t}$ is the flow map generated by $u_{t}$, and the integral is taken along the characteristic path from $t=0$ to $t=1$.
In high dimensions, explicitly computing the divergence $\nabla\cdot u_{t}=\mathrm{tr}(\partial u_{t}/\partial x)$ is computationally expensive ($\mathcal{O}(d^{2})$). To ensure numerical stability and scalability, we employ the Hutchinson trace estimator (hutchinson1989stochastic), which approximates the divergence using random noise vectors $\epsilon\sim\mathcal{N}(0,\mathbf{I})$:
$\nabla\cdot u_{t}(x)=\mathbb{E}_{\epsilon}\left[\epsilon^{\top}\frac{\partial u_{t}(x)}{\partial x}\epsilon\right].$ (35)
This reduces the complexity to $\mathcal{O}(d)$ using vector-Jacobian products. In our experiments, we solve the combined ODE for the particle trajectory and the log-density change using the torchdiffeq library with an adaptive step-size solver (dopri5) and tolerances of $10^{-5}$.
Appendix J Supplementary figures
@@ Appendix I Negative Log-Likelihood Computation
Figure A5: Ablation studies corresponding to Fig. 4. a Effect of doubling the bandwidth $k_{bw}$ on the CDC field at a given value ($k=128$). Minimal effect observed. b Effect of doubling the number of nearest neighbours $k$ on the CDC field at a given value ($k_{bw}=8$). The CDC field is less sensitive to changes in $k$ for higher $k$ values. c Neighbourhood of a point in a dense (left) and sparse (right) manifold region. Sparse regions can lead to shortcut connections that do not respect the manifold geodesic distance. The kernel density estimate assigns large mass to these points, biasing the CDC estimate. d Influence of the choice of $k$ on the overall generalisation and memorisation performance. CDC-FM provides gains over FM for all choices of $k$. e Generalisation and memorisation at the final epoch (100k, mean and std over 3 seeds). The overall performance gain (although improved compared to FM for all choices of $k$) is in a tradeoff between generalisation and memorisation. Small $k$ leads to better tangent space estimates and thus better generalisation, but yields less strong regularisation. Large $k$ leads to worse tangent space estimates, so generalisation drops, but overall reduces memorisation.